OpenStudy (anonymous):

Which of the following is not a step used when graphing the equation 2x + y = 14 using the x- and y- intercepts method? 1.Count down 2 and right 1 from the y-intercept. 2.Plot the y-intercept (0, 14). 3. Plot the intercept (7, 0). 4.Draw a line through the two points.

OpenStudy (anonymous):

To find the y intercept turn the x into zero and solve for y This is 2. To find the x intercept turn y into zero and solve for x This is 3. To find the slope use your points (7,0)(0,14) y2 - y1/x2 - x1 = 14 - 0/0 - 7 = 14/-7 = - 2/1 This is 1. So, your answer is the 4th one :)